# Blind Free Deconvolution over one-parameter sparse families via eigenmatrix

### v1

@@ 1. Introduction
This note considers two blind free deconvolution problems over sparse spectral measures. In the additive setting, $A_{1}$ and $A_{2}$ are both $N\times N$ symmetric matrices with sparse spectral measures ${\mu_{A_{1}}}$ and ${\mu_{A_{2}}}$. Given the spectral measure ${\mu_{C}}$ of $C=A_{1}+QA_{2}Q^{\mathsf{T}}$, where $Q$ is sampled from the Haar measure over orthogonal matrices, the task is to recover ${\mu_{A_{1}}}$ and ${\mu_{A_{2}}}$. This problem can also be extended to the free sum of more than two members.
In the multiplicative setting, $A_{1}$ and $A_{2}$ are both $N\times N$ symmetric positive definite matrices with sparse spectral measures. Given ${\mu_{C}}$ of $C=\sqrt{A_{1}}QA_{2}Q^{\mathsf{T}}\sqrt{A_{1}}$, where $Q$ is again from the Haar measure over orthogonal matrices, the task is to recover the spectral measures ${\mu_{A_{1}}}$ and ${\mu_{A_{2}}}$. This problem can also be extended to the free product of more than two members.
These problems are referred to as blind deconvolution since neither ${\mu_{A_{1}}}$ nor ${\mu_{A_{2}}}$ is known. Blind deconvolutions are typically ill-posed in general. In order to work with well-posed problems, we restrict to sparse spectral measures. Even with the sparsity constraint, the problem is still difficult since the composition from the spectral measures of $A_{1}$ and $A_{2}$ to that of $C$ is nonlinear, hence rendering these two problems as nonlinear inverse problems.
@@
There is a significant body of work on spectral estimation, including linear shrinkage [ledoit2004well], optimization approach using the Marcenko-Pastur equation [el2008spectrum], nonlinear shrinkage methods [ledoit2011eigenvectors, ledoit2012nonlinear, ledoit2015spectrum, ledoit2020analytical, gavish2014optimal, bun2016rotational], moment-based method [kong2017spectrum], and the subordination method [arizmendi2020subordination]. In most of the previous work, the estimation task is not of the blind deconvolution type, as one of the convolution components (such as the noise level or the dimension-to-sample size ratio) is assumed to be known. On the other hand, relatively few works have focused on blind free deconvolution. Recently, the work in [ying2025sparse] considers the spectral estimation with an unknown noise level, which can be viewed as a relatively simpler version of blind deconvolution.
@@
The main technical tool is the recently proposed eigenmatrix method [ying2024eigenmatrix, ying2025multidimensional] for solving very general unstructured sparse recovery problems. This level of generality will play a key role in addressing the problems here.
There are two key ideas. First, the nonlinear inverse problems can be turned into linear ones by leveraging the R-transform for free addition and the S-transform for free product [voiculescu1986addition, voiculescu1987multiplication, mingo2017free, potters2020first]. Once this is done, we extend the eigenmatrix method to the parametric domain case to solve these linear sparse recovery problems.
@@ 2. Eigenmatrix
is the unknown sparse signal, where $\{x_{k}\}_{1\leq k\leq{n}}$ are the spike locations and $\{w_{k}\}_{1\leq k\leq{n}}$ are the spike weights. The observed data is defined via the function
@@
Let $\{z_{j}\}$ be a set of ${n_{z}}$ unstructured samples, and we are only given the noisy observations $u_{j}\approx u(z_{j})$. The task is to recover the spikes $\{x_{k}\}$ and weights $\{w_{k}\}$ from $\{u_{j}\}$.
@@
To build $M$, we pick a Chebyshev grid $\{c_{t}\}_{1\leq t\leq{n_{c}}}$ of size ${n_{c}}$ on the interval $X$. Here ${n_{c}}$ is sufficiently large yet the vectors $\{\hat{\mathbf{b}}_{c_{t}}\}$ are numerically linearly independent. This condition is then enforced on this Chebyshev grid, i.e.,
@@
where the pseudoinverse $\hat{B}^{+}$ is computed by thresholding the singular values of $\hat{B}$. In practice, the thresholding value is chosen so that a small constant bounds the norm of $M$.
The rest of the method follows, for example, the ESPRIT algorithm [roy1989esprit] (or the Prony’s method [prony1795essai]). Define the vector
@@
with ${n_{l}}>{n}$, obtained from applying $M$ repeatitively. Since $\mathbf{u}\approx\sum_{k}\mathbf{b}_{x_{k}}w_{k}$ and $M\mathbf{b}_{x}\approx x\mathbf{b}_{x}$,
@@ 3. Classical deconvolution
This section uses classical deconvolution as a simple example to illustrate the main ideas behind the approach. Let the interval $X$ be the parametric domain of the 1-parameter family of distributions $\rho_{x}(y)$ over $\mathbb{R}$. Each distribution $\alpha_{k}$ is associated with the parameter $x_{k}$, i.e., $\alpha_{k}(y)=\rho_{x_{k}}(y)$. Given $\gamma=\alpha_{1}*\ldots*\alpha_{n}$, the task is to recover the parameters $x_{1},\ldots,x_{n}$, and hence $\alpha_{1},\ldots,\alpha_{n}$.
@@
The task is now to recover $x_{1},\ldots,x_{n}$ from $\log\hat{\gamma}(\xi)$. This is a linear sparse inverse problem with the kernel $G(\xi,x)=\log\hat{\rho}_{x}(\xi)$.
The second step solves this problem using eigenmatrix. We pick a set of samples $\xi_{j}\in\mathbb{R}$ and define for each $x\in X$ the vector
@@
and form the $T$ matrix to recover $x_{1},\ldots,x_{n}$, and equivalently $\alpha_{1},\ldots,\alpha_{n}$.
@@ 4. Free additive deconvolution
For each $A_{k}$, $\mu_{A_{k}}=\rho_{x_{k}}$, and its R-transform is denoted as $r_{\rho_{x_{k}}}(g)$. In the large dimension limit, $r_{\mu_{C}}(g)=r_{\mu_{A_{1}}}(g)+\ldots r_{\mu_{A_{n}}}(g)$, [voiculescu1986addition] or equivalently,
@@
The task is, given $r_{\mu_{C}}(g)$, to recover $x_{1},\ldots,x_{n}$. This is now a linear sparse inverse problem with kernel given by $G(g,x)=r_{\rho_{x}}(g)$.
The second step solves this problem using eigenmatrix. Assume for now that for any $\mu$, we are able to compute $r_{\mu}(g)$. We choose a set of samples $\{g_{j}\}$ on a small circle around the origin. For each $x\in X$, define the vector
@@
Below, we present a few numerical examples. In each example, we choose a 1-parameter family to have the first moment equal to zero. The reason is that a non-zero first moment can be absorbed in any $\mu_{A_{k}}$ with a uniform spectral shift.
@@
  • The 1-parameter family is $\rho_{x}=\frac{1}{2}\delta_{-x}+\frac{1}{2}\delta_{x}$ for $x\in[0.4,1]$.
@@
$\mathbf{b}_{x}:=[r_{\rho_{x}}(g_{j})]_{j}$ is analytic in $x$. Though the condition number of $\hat{B}$ is quite large, the construction of $M$ is still quite accurate. Figure 4.1 summarizes the result. The first plot is the histogram of ${\mu_{C}}$. The second plot gives the exact (red) and reconstructed (blue) parameters $\{x_{k}\}$.
@@
  • The 1-parameter family is $\rho_{x}=\frac{2}{3}\delta_{-x/2}+\frac{1}{3}\delta_{x}$ for $x\in[0.4,1]$.
  • The spectral measure of $A_{1},A_{2}$ correspond to $(x_{1},x_{2})=(0.5,0.9)$.
@@
$\mathbf{b}_{x}:=[r_{\rho_{x}}(g_{j})]_{j}$ is analytic in $x$. Though the condition number of $\hat{B}$ is quite large, the construction of $M$ is still reasonably accurate. Figure 4.2 summarizes the result. The first plot is the histogram of ${\mu_{C}}$. The second plot gives the exact (red) and reconstructed (blue) parameters $\{x_{k}\}$.
@@ 5. Free multiplicative deconvolution
Let the interval $X$ be the parametric domain of the 1-parameter family $\rho_{x}$. For each $k$, the spectral measure $\mu_{A_{k}}=\rho_{x_{k}}$ for unknown $x_{k}$. Given the spectral measure ${\mu_{C}}$ of the product $C=\sqrt{A_{1}}Q_{2}\sqrt{A_{2}}\dots Q_{n}A_{n}Q_{n}^{\mathsf{T}}\ddots\sqrt{A_{2}}Q_{2}^{\mathsf{T}}\sqrt{A_{1}}$, the task is to recover $\mu_{A_{1}},\ldots,\mu_{A_{n}}$ or equivalently $x_{1},\ldots,x_{n}$.
@@
The task is, given $\log s_{\mu_{C}}(t)$, to recover $x_{1},\ldots,x_{n}$. This is a linear sparse inverse problem with a kernel $G(t,x)=\log s_{\rho_{x}}(t)$.
The second step solves this problem using eigenmatrix. Assume for now that for any $\mu$, we are able to compute $s_{\mu}(t)$. Choose a set of samples $\{t_{j}\}$ on a small circle around the origin. For each $x\in X$, define the vector
@@
Given $t$, one way for computing $s$ is to recover $z$ first, via solving the equation $t+1=zg_{\mu}(z)$. This is again done with the help of a Newton iteration. Let $z_{i}$ be the current approximation. To find the update $\epsilon$, we write
@@
  • The 1-parameter family is $\rho_{x}=\frac{1}{2}\delta_{2/(1+x)}+\frac{1}{2}\delta_{2x/(1+x)}$ for $x\in[1.4,3]$.
@@
Here $\mathbf{b}_{x}:=[\log s_{\rho_{x}}(t_{j})]_{j}$ is analytic in $x$. Though the condition number of $\hat{B}$ is quite large, the construction of $M$ is still quite accurate. Figure 5.1 summarizes the result. The first plot is the histogram of ${\mu_{C}}$. The second plot gives the exact (red) and reconstructed (blue) parameters $\{x_{k}\}$.
@@
  • The 1-parameter family is $\rho_{x}=\frac{2}{3}\delta_{3/(2+x)}+\frac{1}{3}\delta_{3x/(2+x)}$ for $x\in[1.4,3]$.
@@
$\mathbf{b}_{x}:=[\log s_{\rho_{x}}(t_{j})]_{j}$ again is analytic in $x$. Though the condition number of $\hat{B}$ is quite large, the construction of $M$ is reasonably accurate. Figure 5.2 summarizes the result. The first plot is the histogram of ${\mu_{C}}$. The second plot gives the exact (red) and reconstructed (blue) parameters $\{x_{k}\}$.
@@ 6. Discussion
This note is a preliminary exploration of blind free deconvolution problems. Below are several clear directions for improvements and future work.
@@
Third, the eigenmatrix method can be viewed as a non-Hermitian quantization of the spectral variety $X$. The vectors $\mathbf{b}_{x}$ serve as the approximate non-orthogonal eigenvectors. Once the matching condition $M\mathbf{b}_{x}\approx x\mathbf{b}_{x}$ is satisfied on a sufficiently dense Chebyshev grid, the whole spectral variety $X$ resides in the pseudo-spectrum of $M$. The original eigenmatrix method was proposed for simple $X$ (either an interval or a domain in the complex plane). The work discussed here is a first step towards addressing a more general spectral variety, even though we only work with a curve in it by considering only a one-parametric family.

### v2

@@ 1. Introduction
This note considers two blind free deconvolution problems over sparse spectral measures from one-parameter families. In the additive setting, $A_{1}$ and $A_{2}$ are both $N\times N$ symmetric matrices with spectral measures ${\mu_{A_{1}}}$ and ${\mu_{A_{2}}}$. Given the spectral measure ${\mu_{C}}$ of $C=A_{1}+QA_{2}Q^{\mathsf{T}}$, where $Q$ is sampled from the Haar measure over orthogonal matrices, the task is to recover ${\mu_{A_{1}}}$ and ${\mu_{A_{2}}}$. This problem can also be extended to the free sum of more than two terms: given ${\mu_{C}}$ of $C=A_{1}+Q_{2}A_{2}Q_{2}^{\mathsf{T}}+\ldots+Q_{n}A_{n}Q_{n}^{\mathsf{T}}$, the task is to recover the spectral measures $\mu_{A_{1}},\ldots,\mu_{A_{n}}$.
In the multiplicative setting, $A_{1}$ and $A_{2}$ are both $N\times N$ symmetric positive definite matrices with sparse spectral measures. Given ${\mu_{C}}$ of $C=\sqrt{A_{1}}QA_{2}Q^{\mathsf{T}}\sqrt{A_{1}}$, where $Q$ is again from the Haar measure over orthogonal matrices, the task is to recover the spectral measures ${\mu_{A_{1}}}$ and ${\mu_{A_{2}}}$. This problem can also be extended to the free product of more than two terms: given ${\mu_{C}}$ of $C=\sqrt{A_{1}}Q_{2}\sqrt{A_{2}}\ldots Q_{n}A_{n}Q_{n}^{\mathsf{T}}\ldots\sqrt{A_{2}}Q_{2}^{\mathsf{T}}\sqrt{A_{1}}$, the task is to recover the spectral measures $\mu_{A_{1}},\ldots,\mu_{A_{n}}$.
These problems are referred to as blind free deconvolution since neither ${\mu_{A_{1}}}$ nor ${\mu_{A_{2}}}$ is known. Blind free deconvolutions for general spectral measures are ill-posed. For example, the sum of two semicircle laws is a semicircle law with the sum of their widths, and therefore, it is not possible to recover the individual inputs. In this note, we restrict to sparse spectral measures. Even with the sparsity constraint, the problem remains difficult since the composition of the spectral measures of $A_{1}$ and $A_{2}$ into that of $C$ is nonlinear, thereby rendering these two problems nonlinear inverse problems.
@@
There is a significant body of work on spectral estimation, including the linear shrinkage [ledoit2004well], the optimization approach using the Marcenko-Pastur equation [el2008spectrum], the nonlinear shrinkage methods [ledoit2011eigenvectors, ledoit2012nonlinear, ledoit2015spectrum, ledoit2020analytical, gavish2014optimal, bun2016rotational], the moment-based method [kong2017spectrum], and the subordination method [arizmendi2020subordination]. In most of the previous work, the estimation task is not of the blind deconvolution type, as one of the convolution components (such as the noise level or the dimension-to-sample size ratio) is assumed to be known. Relatively few works have focused on blind free deconvolution. Recently, the work in [ying2025sparse] considers the spectral estimation with an unknown noise level, which can be viewed as a relatively simpler version of blind deconvolution.
@@
The primary technical tool is the recently proposed eigenmatrix method [ying2024eigenmatrix, ying2025multidimensional] for solving general, yet unstructured, sparse recovery problems. This level of generality will play a key role in addressing the problems here.
There are two key ideas. First, the nonlinear inverse problems can be turned into linear ones by leveraging the R-transform for free addition and the S-transform for free product [voiculescu1986addition, voiculescu1987multiplication, mingo2017free, potters2020first]. Once this is done, we extend the eigenmatrix method to the parametric case to solve the corresponding linear sparse recovery problems.
@@ 2. Eigenmatrix
is the unknown sparse signal, where $\{x_{k}\}_{1\leq k\leq{n}}$ are the spike locations, $\{w_{k}\}_{1\leq k\leq{n}}$ are the spike weights, and $\delta(\cdot)$ is the Dirac delta function at the origin. The observed data is defined via the function
@@
Let $\{z_{j}\}$ be a set of ${n_{z}}$ unstructured samples and $u_{j}\approx u(z_{j})$ be the noisy observations of $u(z)$ at the chosen samples. The task is to recover the spikes $\{x_{k}\}$ and weights $\{w_{k}\}$ from $\{u_{j}\}$.
@@
To build $M$, we pick a Chebyshev grid $\{c_{t}\}_{1\leq t\leq{n_{c}}}$ of size ${n_{c}}$ [trefethen2019approximation] of the interval $X$. Here ${n_{c}}$ is sufficiently large yet the vectors $\{\hat{\mathbf{b}}_{c_{t}}\}$ are numerically linearly independent. This condition is then enforced on this Chebyshev grid, i.e.,
@@
where the pseudoinverse $\hat{B}^{+}$ is computed by thresholding the singular values of $\hat{B}$. In practice, the thresholding value is chosen so that a small constant bounds the operator norm of $M$.
The rest of the method follows, for example, the ESPRIT algorithm [roy1989esprit] (or the Prony method [prony1795essai]). Define the vector
@@
with ${n_{l}}>{n}$, obtained from applying $M$ repetitively. Since $\mathbf{u}\approx\sum_{k}\mathbf{b}_{x_{k}}w_{k}$ and $M\mathbf{b}_{x}\approx x\mathbf{b}_{x}$,
@@ 3. Classical deconvolution
This section illustrates the main ideas using the classical deconvolution as an example, which is a new application of the eigenmatrix method. Let the interval $X$ be the parametric domain of the 1-parameter family of distributions $\rho_{x}(y)$ over $\mathbb{R}$. Each distribution $\alpha_{k}$ is associated with the parameter $x_{k}$, i.e., $\alpha_{k}(y)=\rho_{x_{k}}(y)$. Given $\gamma=\alpha_{1}*\ldots*\alpha_{n}$, the task is to recover the parameters $x_{1},\ldots,x_{n}$, and hence $\alpha_{1},\ldots,\alpha_{n}$.
@@
The task is now to recover $x_{1},\ldots,x_{n}$ from $\log\hat{\gamma}(\xi)$. This is a linear sparse inverse problem with the kernel $G(\xi,x)=\log\hat{\rho}_{x}(\xi)$ and the weights $w_{k}$ equal to $1$.
The second step solves this problem using the eigenmatrix method. We pick a set of samples $\xi_{j}\in\mathbb{R}$ and define for each $x\in X$ the vector
@@
and form the $T$ matrix to recover $x_{1},\ldots,x_{n}$ and equivalently $\alpha_{1},\ldots,\alpha_{n}$.
@@ 4. Free additive deconvolution
For each $A_{k}$, $\mu_{A_{k}}=\rho_{x_{k}}$, and its R-transform is denoted as $r_{\rho_{x_{k}}}(g)$. In the large dimension limit, $r_{\mu_{C}}(g)=r_{\mu_{A_{1}}}(g)+\ldots+r_{\mu_{A_{n}}}(g)$, [voiculescu1986addition] or equivalently,
@@
The task is, given $r_{\mu_{C}}(g)$, to recover $x_{1},\ldots,x_{n}$. This is now a linear sparse inverse problem with the kernel given by $G(g,x)=r_{\rho_{x}}(g)$ and the weights $w_{k}$ all equal to $1$.
The second step solves this problem using the eigenmatrix method. Assume for now that for any $\mu$, we are able to compute $r_{\mu}(g)$. We choose a set of samples $\{g_{j}\}$ on a small circle around the origin. For each $x\in X$, define the vector
@@
Below, we present a few numerical examples. In each example, we select a 1-parameter family with the first moment equal to zero. The reason is that a non-zero first moment can be absorbed in any $\mu_{A_{k}}$ with a uniform spectral shift.
@@
  • The 1-parameter family is $\rho_{x}=\frac{2}{3}\delta_{-x/2}+\frac{1}{3}\delta_{x}$ for $x\in X=[0.4,1]$.
  • The spectral measure of $A_{1},A_{2}$ correspond to $(x_{1},x_{2})=(0.5,0.9)$.
@@
Here $\mathbf{b}_{x}:=[r_{\rho_{x}}(g_{j})]_{j}$ is analytic in $x$. Though the condition number of $\hat{B}$ is quite large, the construction of $M$ is quite accurate. Figure 4.1 summarizes the result. The first plot is the histogram of ${\mu_{C}}$. The second plot gives the exact (red) and reconstructed (blue) parameters $\{x_{k}\}$.
@@
  • The 1-parameter family is $\rho_{x}=\frac{1}{2}\delta_{-x}+\frac{1}{2}\delta_{x}$ for $x\in X=[0.4,1]$.
@@
Here $\mathbf{b}_{x}:=[r_{\rho_{x}}(g_{j})]_{j}$ is analytic in $x$. Though the condition number of $\hat{B}$ is quite large, the construction of $M$ is still quite close. Figure 4.2 summarizes the result. The first plot is the histogram of ${\mu_{C}}$. The second plot gives the exact (red) and reconstructed (blue) parameters $\{x_{k}\}$.
@@ 5. Free multiplicative deconvolution
Let the interval $X$ be the parametric domain of the 1-parameter family $\rho_{x}$. For each $k$, the spectral measure $\mu_{A_{k}}=\rho_{x_{k}}$ for unknown $x_{k}$. Given the spectral measure ${\mu_{C}}$ of the product $C=\sqrt{A_{1}}Q_{2}\sqrt{A_{2}}\ldots Q_{n}A_{n}Q_{n}^{\mathsf{T}}\ldots\sqrt{A_{2}}Q_{2}^{\mathsf{T}}\sqrt{A_{1}}$, the task is to recover $\mu_{A_{1}},\ldots,\mu_{A_{n}}$ or equivalently $x_{1},\ldots,x_{n}$.
@@
The task is, given $\log s_{\mu_{C}}(t)$, to recover $x_{1},\ldots,x_{n}$. This is a linear sparse inverse problem with the kernel $G(t,x)=\log s_{\rho_{x}}(t)$ and all weights $w_{k}$ equal to 1.
The second step solves this problem using the eigenmatrix method. Assume for now that for any $\mu$, we are able to compute $s_{\mu}(t)$. Choose a set of samples $\{t_{j}\}$ on a small circle around the origin. For each $x\in X$, define the vector
@@
Given $t$, one way to compute $s$ is to recover $z$ first, via solving the equation $t+1=zg_{\mu}(z)$. This is again done with the help of a Newton iteration. Let $z_{i}$ be the current approximation. To find the update $\epsilon$, we write
@@
  • The 1-parameter family is $\rho_{x}=\frac{2}{3}\delta_{3/(2+x)}+\frac{1}{3}\delta_{3x/(2+x)}$ for $x\in X=[1.4,3]$.
@@
$\mathbf{b}_{x}:=[\log s_{\rho_{x}}(t_{j})]_{j}$ again is analytic in $x$. Though the condition number of $\hat{B}$ is quite large, the construction of $M$ is quite accurate. Figure 5.1 summarizes the result. The first plot is the histogram of ${\mu_{C}}$. The second plot gives the exact (red) and reconstructed (blue) parameters $\{x_{k}\}$.
@@
  • The 1-parameter family is $\rho_{x}=\frac{1}{2}\delta_{2/(1+x)}+\frac{1}{2}\delta_{2x/(1+x)}$ for $x\in X=[1.4,3]$.
@@
Here $\mathbf{b}_{x}:=[\log s_{\rho_{x}}(t_{j})]_{j}$ is analytic in $x$. Though the condition number of $\hat{B}$ is quite large, the construction of $M$ is still quite close. Figure 5.2 summarizes the result. The first plot is the histogram of ${\mu_{C}}$. The second plot gives the exact (red) and reconstructed (blue) parameters $\{x_{k}\}$.
@@ 6. Discussion
This note is a preliminary exploration of blind free deconvolution problems. Below are several directions for improvements and future work.
@@
Third, the eigenmatrix method can be viewed as a non-Hermitian quantization of the spectral variety $X$. The vectors $\mathbf{b}_{x}$ serve as the approximate non-orthogonal eigenvectors. Once the matching condition $M\mathbf{b}_{x}\approx x\mathbf{b}_{x}$ is satisfied on a sufficiently dense Chebyshev grid, the whole spectral variety $X$ resides in the pseudo-spectrum of $M$. The original eigenmatrix method was proposed for simple $X$ (either an interval or a domain in the complex plane). The work here is a first step towards addressing a more general spectral variety, even though we consider only a curve within it by focusing on a one-parameter family.